Maple Syrup worth $6 a gallon and corn syrup worth 80 cents a gallon are used to make a mixture worth $2.36 a gallon. How many gallons of each kind of syrup are needed to make 50 gallons of the mixture?
I have to find out how much of each is in the mixtured solution that costs $2.36 and then multiply those numbers by 50, right? I think what's confusing me most is the 2 variables.
Thank you!

Answer by checkley71(8403)

600x+80(50-x)=236x50 or 600x+4000-80x=11,800 or 520x=7,800 or x=7,800/520 or
x=15 so 600x15+80(50-15)=11,800 or 9,000+4,000-1,200=11,800 or
13,000-1,200=11,800 or
15 gallons of $6.00 maple syrup plus 35 gallons of $0.80 corn syrup to make
50 gallons of $ 2.36 mix.